OpenStudy (anonymous):

Let ln2=a, ln3=b, and ln5=c. Express the following logarithms in terms of abc . ln800= ln0.096= a+a+b-3c ln0.027= 3b-(3a+3c) log2 27000= (3a+3b+3c)/a Cant find ln800

OpenStudy (ranga):

Find the prime factors of 800 first.

OpenStudy (raden):

1# setting 800 = 2^5 * 5^2 so, ln800 = ln(2^5 * 5^2) = ln(2^5) + ln(5^2) = 5ln2 + 2ln5 = 5a + 2c

OpenStudy (anonymous):

Yeah I just got it as you guys replied. Thanks though!